An IQ test is designed so that the mean is 100 and the standard deviation is 23 for the population of normal adults. Find the sample size necessary to estimate the mean IQ score of statistics students such that it can be said with 95​% confidence that the sample mean is within 6 IQ points of the true mean. Assume that sigma σ equals = 23 and determine the required sample size using technology. Then determine if this is a reasonable sample size for a real world calculation. The required sample size is nothing . ​(Round up to the nearest​ integer.)
Answer #1

Solution :

Given that,

Population standard deviation = = 23

Margin of error = E = 6

At 95% confidence level

= 1-0.95% =1-0.95 =0.05

/2 =0.05/ 2= 0.025

Z/2 = Z0.025 = 1.960

Z/2 = 1.960  

sample size = n = [Z/2* / E] 2

n = [1.960 *23 /6 ]2

n = 56

Sample size = n = 56